OpenStudy (anonymous):

Find the volume of the square pyramid shown. Round to the nearest tenth if necessary.

OpenStudy (anonymous):

5,184 ft3 1,728 ft3 192 ft3 2,592 ft3

OpenStudy (callisto):

 \[h = \sqrt{15^2 - 12^2} = 9\] Volume = (1/3)(24x24x9) = ?